OpenStudy (anonymous):

calculate the moment of inertia of a circular solid disk of radius R from which another disk from the top most part of radius R/3 has been cut out

OpenStudy (anonymous):



OpenStudy (anonymous):

The axis about which MI is to be found passes through center???

OpenStudy (anonymous):

yes through the center

OpenStudy (anonymous):

Do u know the answer???

OpenStudy (anonymous):

yes its 4MR^2 where M is the mass of the solid disk

OpenStudy (anonymous):

The answer should be (4MR^2)/9

OpenStudy (anonymous):

ohh sorry its mass is 9M

OpenStudy (anonymous):

please answer I am sorry for the mistake

OpenStudy (anonymous):

yup then my answer is correct See, MI = (MR^2)/2 (normal disc) \[Let\ \sigma\ \]be the Mass per unit area So, \[\sigma = M/\pi r ^{2}\] So the mass of disk given becomes 8/9(M). So acc. to formula \[[8/9(M) \times \pi r ^{2}]/2 = 4/9Mr ^{2}\]

OpenStudy (anonymous):

Your mass M=9M

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

from did (8/9)M come from

OpenStudy (anonymous):

ur area is pi(r^2) - pi{(r/3)^2} = 8/9(pi)(r^2)

OpenStudy (anonymous):

whats the formulae u used for last step

OpenStudy (anonymous):

MI = (MR^2)/2 (for a disc)

OpenStudy (anonymous):

why did you multiply pi

OpenStudy (anonymous):

ok thanks